OpenStudy (anonymous):

The rate of decay in the mass, M, of a radioactive substance is given by the differential equation dM dt equals negative 1 times k times M, where k is a positive constant. If the initial mass was 150g, then find the expression for the mass, M, at any time t.

OpenStudy (anonymous):

https://gyazo.com/8d62d90df17bfc829809548f15163bd1

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

Can you confirm my answer?

OpenStudy (anonymous):

@baru Can you confirm for me please?

OpenStudy (anonymous):

@jim_thompson5910 will you be my hero?

jimthompson5910 (jim_thompson5910):

If \(\Large M = 150e^{-kt}\) then what is \(\Large \frac{dM}{dt}\) equal to?

OpenStudy (anonymous):

-kM

OpenStudy (anonymous):

@jim_thompson5910 when I integrated I got M=e^-kt

OpenStudy (anonymous):

I wasn't sure what to do with the initial mass, so using the choices I multiplied it.

OpenStudy (anonymous):

Would you like to see my work?

jimthompson5910 (jim_thompson5910):

sure, go ahead and post your work

OpenStudy (anonymous):

Please ignore the 100-60 crap, that was for another question.

jimthompson5910 (jim_thompson5910):

You should have M in absolute values. And don't forget the +C \[\Large \ln(|M|) = -kt+C\] \[\Large |M| = e^{-kt+C}\] \[\Large |M| = e^{C}*e^{-kt}\]

OpenStudy (anonymous):

Why in absolute values?

jimthompson5910 (jim_thompson5910):

because the domain of ln(x) is x > 0 the absolute values ensure that the argument is not negative

OpenStudy (anonymous):

Ahh ok.

OpenStudy (anonymous):

Would e^c be equal to 150?

jimthompson5910 (jim_thompson5910):

now you'll use the information that 150 is the initial mass so M = 150 when t = 0 \[\Large |M| = e^{C}*e^{-kt}\] \[\Large |150| = e^{C}*e^{-k*0}\] \[\Large |150| = e^{C}*e^{0}\] \[\Large |150| = e^{C}*1\] \[\Large |150| = e^{C}\] \[\Large 150 = e^{C}\] \[\Large e^{C} = 150\] you are correct

OpenStudy (anonymous):

Thank you good sir.

jimthompson5910 (jim_thompson5910):

Now we make the claim that \[\Large M = 150e^{-kt}\] here is how to check that claim If \[\Large M = 150e^{-kt}\] then \[\Large \frac{dM}{dt} = -150ke^{-kt}\] -------------------------------- Start with the initial differential equation and perform substitutions \[\Large \frac{dM}{dt} = -k*M\] \[\Large -150ke^{-kt} = -k*150e^{-kt}\] \[\Large -150ke^{-kt} = -150ke^{-kt} \ \ \ \color{green}{\checkmark}\] So the solution has been confirmed

jimthompson5910 (jim_thompson5910):

Of course, the more general solution is \[\Large M = Ce^{-kt}\] but we don't need to worry about that

OpenStudy (anonymous):

Got it. Thanks